The price of a BMW is $33,000 in the United States and e28,000 in Germany. Where...

  1. The price of a BMW is $33,000 in the United States and e28,000 in Germany.
  1. Where would you buy? Why?
  1. Assume one U.S. dollar buys e0.83. Is there any arbitrage opportunity? Where would you buy? Why?
  1. Compute the no-arbitrage e$ exchange rate. At this exchange rate, where would you buy?
Answer #1

a. As such untill we don't know the exchanged rate of Euro to dollar we are not able to find the best option let's take current exchange rate : Euro is equal to 1.10 USD. 1€=1$.

So if we buy it from Germany price will be = 28000*1.10 = 30800$. And if we buy from USA it will cost us 33000$ so we should buy it from Germany only.

b. 1$ = 0.83€

Then if we buy from Germany = 28000€

If we buy from USA = 33000$*0.83€ = 27390€

Is there arbitrage opportunity??

Yes to establish arbitrage we will buy it from America and sell it into the Germany and we can make arbitrage profit of = 28000-27390 = 610€

In this situation I should buy it from USA.

c. no arbitrage exchange rate :-

33000$= 28000€

1$ = 28000/33000 €

1$ = 0.85€

At this exchange rate buyer will be indifferent about the country choice because the cost of the car will be the same whether we buy it from Germany or USA. It will cost us the same.

i hope my efforts will be fruitful to you....

Waiting for your feedback...?